# FinCoT: Grounding Chain-of-Thought in Expert Financial Reasoning

### v2

@@ Abstract
This paper presents FinCoT, a structured chain-of-thought (CoT) prompting framework that embeds domain-specific expert financial reasoning blueprints to guide large language models’ behaviors. We identify three main prompting styles in financial NLP (FinNLP): (1) standard prompting (zero-shot), (2) unstructured CoT (free-form reasoning), and (3) structured CoT (with explicitly structured reasoning steps). Prior work has mainly focused on the first two, while structured CoT remains underexplored and lacks domain expertise incorporation. Therefore, we evaluate all three prompting approaches across ten CFA-style financial domains and introduce FinCoT as the first structured finance-specific prompting approach incorporating blueprints from domain experts. FinCoT improves the accuracy of a general-purpose model, Qwen3-8B-Base, from 63.2% to 80.5%, and boosts Fin-R1 (7B), a finance-specific model, from 65.7% to 75.7%, while reducing output length by up to 8.9$\times$ and 1.16$\times$ compared to structured CoT methods, respectively. We find that FinCoT proves most effective for models lacking financial post-training. Our findings show that FinCoT does not only improve performance and reduce inference costs but also yields more interpretable and expert-aligned reasoning traces.
@@ 4 Experimental Setup
We report accuracy as the metric, defined as the percentage of questions where the model’s prediction matches the ground truth. To assess response efficiency, we also report the average output length (in tokens) across questions. For statistical significance, we use a paired bootstrap test (Efron and Tibshirani, 1994) with $B{=}10k$ resamples over binary correctness scores, reporting the mean difference ($\Delta$), 95% confidence interval, and $p$-value. Additionally, we measure the proportion of financial domains where a method improves accuracy by at least 1% over SP and compute the average domain-wise accuracy gain.
@@ 5.1 Financial Reasoning Performance
Table 1: Comparison of accuracy (%) of prompting techniques. ‘FinCoT’ simultaneously applies expert reasoning blueprints from all CFA domains, while each ‘(DomainName)’ (e.g., ‘Economics’) row applies domain-specific blueprints individually. ($\uparrow$/$\downarrow$) Denote accuracy improvement or decline relative to the SP baseline, colored green for ($\uparrow$) and red for ($\downarrow$). Bold values highlight the best-performing prompt variant for each model. (*) Indicates that the accuracy improvement among the model-level prompt variants is statistically significant ($p<0.05$) based on paired bootstrap testing; domain-specific rows are not tested for significance.
@@
Table˜1 reports zero-shot accuracies for four prompting strategies (SP, UST-CoT, ST-CoT, FinCoT) across our model suite. Under the basic SP prompt, the instruction-tuned Qwen3-8B (Thinker) attains the highest accuracy among general-purpose models (88.18%), while the financial model Fin-o1 (8B) leads its group at 79.65%. These strong baselines clearly highlight the effectiveness of instruction tuning and domain specificity.
@@ 5.2 Efficiency Analysis
For general-purpose models such as Qwen3-8B-Base and Qwen3-8B (Thinker), FinCoT substantially reduces output length while preserving or improving accuracy. On Qwen3-8B-Base, FinCoT improves accuracy from 78.49% (ST-CoT) to 80.52% (+2.03 pp), while reducing average output from 3.42k to 0.38k tokens, an approximately 8.9$\times$ compression. Similarly, on Qwen3-8B (Thinker), FinCoT maintains parity with ST-CoT at 88.18% accuracy, using 1.23k tokens compared to 1.35k, a $\sim$1.1$\times$ reduction. These results support the view that FinCoT’s structured blueprints enable more concise, focused reasoning in general-purpose models. Among financial-specific models such as DianJin-R1-7B and Fin-o1-8B, FinCoT shows minimal or negative improvement and little benefit in output compression. This aligns with the idea that fine-tuned models already internalize domain-specific reasoning. In particular, Fin-o1-8B illustrates that excessive prompt scaffolding may interfere with latent reasoning, reducing effectiveness and leading to what can be described as overthinking that undermines transfer.
@@ 6 Conclusion
We presented FinCoT, a zero-shot prompting framework that embeds expert-curated Mermaid diagrams within structured chain-of-thought scaffolds. By grounding reasoning in domain logic, FinCoT bridges human financial workflows with LLM outputs, without model fine-tuning. While broadly applicable, FinCoT yields strong gains for general models (e.g., Qwen3-8B-Base), but more modest or negative effects for instruction- or finance-specific models (e.g., Qwen-Thinker, Fin-o1, DianJin-R1), where added structure may interfere with learned reasoning.
Relative to SP, FinCoT improves Qwen3-8B-Base by +17.33 pp and Fin-R1 by +10.08 pp ($p<0.001$), outperforming some fine-tuned models. In contrast, instruction-tuned models like Qwen3-8B (Thinker) sometimes favor UST-CoT. Cross-domain results show blueprints from quantitative fields transfer best (+27.3 pp on Qwen3-8B-Base). FinCoT also reduces token output by up to 8×8$\times$8 ×, offering interpretable, efficient prompting for regulated financial applications. Ultimately, FinCoT suggests that with meticulous prompt design, even general-purpose LLMs can approach the reasoning quality of fine-tuned financial experts in complex decision-making tasks.
@@ Acknowledgments
We extend our gratitude to the Hatari Team for generously providing additional GPU resources, which significantly enhanced our computational capacity and enabled the successful execution of large-scale experiments. We also thank Tuntai Boriboonthana (CFA Level-III holder) from InnovestX for his expert contributions in verifying blueprint fidelity and conducting random audits to ensure domain classification consistency in GPT-4o outputs. Their support and insights were invaluable to the development and evaluation of this work.
@@ Appendix A Expert Reasoning Blueprints
Step-4: Determine Numerical vs. Conceptual Approach (C) – Numerical (D1): Gather the relevant data (prices, quantities, rates), apply formulae (e.g., $\text{Elasticity}=\frac{\%\Delta Q_{n}}{\%\Delta P}$, $)\text{GDP}=C+I+G+(X-M)$, and verify assumptions (Investopedia, 2025c). – Conceptual (D2): Construct a narrative explaining cause–effect relationships (e.g., how a monetary-policy change shifts AD or how trade barriers affect capital flows) (Investopedia, 2025b).
@@
300Hours CFA Level 1 Economics Cheat Sheet (300Hours, 2025a), UWorld Finance’s CFA® Economics: Syllabus & Sample Questions (UWorld Finance, 2025a), Kaplan Schweser’s Level I Economics tips (Kaplan Schweser, 2025), Efficient Learnings CFA Economics overview (Efficient Learning, 2025), iPass Finance Exams’ study guide (iPassFinanceExams, 2025), AnalystPrep’s essential review summary (AnalystPrep, 2025), and key Investopedia articles on supply and demand (Investopedia, 2025a), GDP (Investopedia, 2025c), and business cycles (Investopedia, 2025b) as well as Prof. Brian Gordon’s CFA Exam Level 1 Economics video (Gordon, 2020a).
@@
    %Notes and detailed steps
@@
Step-5: Develop Recommendations – Formulate strategies such as adjusting overall duration (shorten if rates likely to rise), implementing barbell or laddered maturity structures, or choosing bullet portfolios to manage reinvestment and rate risk (James, 2025).
@@
    %Step-specific highlights
@@
    %Example labels or notes (optional)
@@ E.1 Accuracy
Table 3: Paired bootstrap significance testing ($B=10{,}000$ samples) for accuracy differences across prompt strategies. $\Delta$ indicates average accuracy difference (in percentage points), with 95% confidence intervals (CI) and $p$-values. A result is considered statistically significant if $p<0.05$.
@@ E.2 Average Output Tokens
Table 4: Paired bootstrap significance testing ($B=10{,}000$ samples) for average output token differences across prompt strategies. $\Delta$ indicates mean difference in output length (in thousands of tokens), with 95% confidence intervals and $p$-values. A result is significant if $p<0.05$.

### v3

@@ Abstract
[…]
@@ 4 Experimental Setup
We report accuracy as the metric, defined as the percentage of questions where the model’s prediction matches the ground truth. To assess response efficiency, we also report the average output length (in tokens) across questions. For statistical significance, we use a paired bootstrap test (Efron and Tibshirani, 1994) with B=10kB{=}10k resamples over binary correctness scores, reporting the mean difference ($\Delta$), 95% confidence interval, and $p$-value. Additionally, we measure the proportion of financial domains where a method improves accuracy by at least 1% over SP and compute the average domain-wise accuracy gain.
@@ 5.1 Financial Reasoning Performance
Table 1: Comparison of accuracy (%) of prompting techniques. ‘FinCoT’ simultaneously applies expert reasoning blueprints from all CFA domains, while each ‘(DomainName)’ (e.g., ‘Economics’) row applies domain-specific blueprints individually. ($\uparrow$$\downarrow$) Denote accuracy improvement or decline relative to the SP baseline, colored green for ($\uparrow$) and red for ($\downarrow$). Bold values highlight the best-performing prompt variant for each model. (*) Indicates that the accuracy improvement among the model-level prompt variants is statistically significant ($p<0.05$) based on paired bootstrap testing; domain-specific rows are not tested for significance.
@@
Table˜1 reports zero-shot accuracies for four prompting strategies (SP, UST-CoT, ST-CoT, FinCoT) across our model suite. Under the basic SP prompt, the instruction-tuned Qwen3-8B (Thinker) attains the highest accuracy among general-purpose models (88.18%), while the financial model Fin-o1-8B leads its group at 79.65%. These strong baselines clearly highlight the effectiveness of instruction tuning and domain specificity.
@@ 5.2 Efficiency Analysis
[…]
@@ 6 Conclusion
We presented FinCoT, a zero-shot prompting framework that embeds expert-curated Mermaid diagrams within structured chain-of-thought scaffolds. By grounding reasoning in domain logic, FinCoT bridges human financial workflows with LLM outputs, without model fine-tuning. While broadly applicable, FinCoT yields strong gains for general models (e.g., Qwen3-8B-Base), but more modest or negative effects for instruction or finance-specific models (e.g., Qwen-Thinker, Fin-o1, DianJin-R1), where added structure may interfere with learned reasoning.
Relative to SP, FinCoT improves Qwen3-8B-Base by +17.33 pp and Fin-R1 by +10.08 pp ($p<0.001$), outperforming some fine-tuned models. In contrast, instruction-tuned models like Qwen3-8B (Thinker) sometimes favor UST-CoT. Cross-domain results show blueprints from quantitative fields transfer best (+27.3 pp on Qwen3-8B-Base). FinCoT also reduces token output by up to $8\times$, offering interpretable, efficient prompting for regulated financial applications. Ultimately, FinCoT suggests that with meticulous prompt design, even general-purpose LLMs can approach the reasoning quality of fine-tuned financial experts in complex decision-making tasks.
@@ Acknowledgments
We extend our gratitude to the Hatari Team for generously providing additional GPU resources, which significantly enhanced our computational capacity and enabled the successful execution of large-scale experiments. We also thank Tuntai Boriboonthana (CFA Charterholder) from InnovestX for his expert contributions in verifying blueprint fidelity and conducting random audits to ensure domain classification consistency in GPT-4o outputs. Their support and insights were invaluable to the development and evaluation of this work.
@@ Appendix A Expert Reasoning Blueprints
Step-4: Determine Numerical vs. Conceptual Approach (C) – Numerical (D1): Gather the relevant data (prices, quantities, rates), apply formulae (e.g., $\text{Elasticity}=\frac{\%\Delta Q_{n}}{\%\Delta P}$, $\text{GDP}=C+I+G+(X-M)$ ), and verify assumptions (Investopedia, 2025c). – Conceptual (D2): Construct a narrative explaining cause–effect relationships (e.g., how a monetary-policy change shifts AD or how trade barriers affect capital flows) (Investopedia, 2025b).
@@
300Hours CFA Level 1 Economics Cheat Sheet (300Hours, 2025a), UWorld Finance’s CFA® Economics: Syllabus & Sample Questions (UWorld Finance, 2025a), Kaplan Schweser’s Level I Economics tips (Kaplan Schweser, 2025), Efficient Learning’s CFA Economics overview (Efficient Learning, 2025), iPass Finance Exams’ study guide (iPassFinanceExams, 2025), AnalystPrep’s essential review summary (AnalystPrep, 2025), and key Investopedia articles on supply and demand (Investopedia, 2025a), GDP (Investopedia, 2025c), and business cycles (Investopedia, 2025b) as well as Prof. Brian Gordon’s CFA Exam Level 1 Economics video (Gordon, 2020a).
@@
    %% Notes and detailed steps
@@
Step-5: Develop Recommendations – Formulate strategies such as adjusting overall duration (shorten if rates are likely to rise), implementing barbell or laddered maturity structures, or choosing bullet portfolios to manage reinvestment and rate risk (James, 2025).
@@
    %% Step-specific highlights
@@
    %% Example labels or notes (optional)
@@ E.1 Accuracy
Table 3: Paired bootstrap significance testing (B=10, [^0] B=10{,}000 samples) for accuracy differences across prompt strategies. $\Delta$ indicates average accuracy difference (in percentage points), with 95% confidence intervals (CI) and $p$-values. A result is considered statistically significant if $p<0.05$.
@@ E.2 Average Output Tokens
Table 4: Paired bootstrap significance testing (B=10, [^0] B=10{,}000 samples) for average output token differences across prompt strategies. $\Delta$ indicates mean difference in output length (in thousands of tokens), with 95% confidence intervals and $p$-values. A result is significant if $p<0.05$.